OpenStudy (anonymous):

Please help ASAP!!! 1. solve the system using matrices { x+ 2y = -1 { -x - 3y = -2 A. [-7 ] [3 ] B. [3 ] [-7] C. [7 ] [-3] D. no solution

OpenStudy (rock_mit182):

\[x-2y=-1 \] \[-x-3y=-2\] now in matrix form: by crAMER'S RULE we have: \[\Delta =\left[\begin{matrix}1 & -2 \\ -1 & -3\end{matrix}\right]\] \[x =\left[\begin{matrix}-1 & -2 \\ -2 & -3\end{matrix}\right]\] \[y=\left[\begin{matrix}1 & -1 \\ -1 & -2\end{matrix}\right]\]